Derive the ODE for the RLC circuit provided using Kirchoff's voltage and current laws. Indicate the order of the equation and what conditions are required for a unique and completely defined solution to existing. You should solve for the voltage across the capacitor as the output, y(t), due to the excitation of the input voltage, x(t), and the initial conditions, y(0) = c0 and y'(0)=c1.

Compute the homogeneous solution to the ordinary differential equation for unknown values of R, L and C. You can keep the solution in exponential form or simplify the expression for cases where the response is under-damped, critically-damped and over-damped.
Answer #1

Page - 3 Homogeneous Sol is y(t) = (co +(c +R (6) x Je ure2 L ²+ & D + į = 0 [Page 2 9 D + 16) - 4x013 (5) 2 for unique and defined solution, we know, + (%) - 4 x0* = 0 = DER so, the| Page 1 Given, Series RLC ekt, 7(0) = co ? initial yo) = c conditions alt; = V مقدوا cyt) fig: 1 solution: - Derivation of OFor the series RLC circuit we can derive the Ordinary Differential Equation by KVL and then can find the unique solution of the Differential equation(making it homogeneous) by standard solution described in the solution.